### v1

@@ 3.6 Continuous delay
Continuous (sometimes called distributed) delays result in Volterra-type integro-differential equations. Driscoll (2010) presents a detailed description of a Chebyshev spectral collocation method for such equations, and we refer the reader to that work for further information. We include an example here both for completeness and to demonstrate that, since both methods are natural extensions to standard Chebyshev spectral methods for ODEs, they can be effortlessly combined to solve problems with a combination of both discrete, proportional, and continuous delay types.
@@
The code in the left panel of Figure 7 demonstrates compact MATLAB code for solving this problem, combining the resampling approach of Section 3.1 with code from (Driscoll, 2010) for discertising the Volterra operator. The computed solution is depicted in the centre panel and geometric convergence as the discretisation size is increased is demonstrated on the right.
@@
Refer to captionRefer to caption
Figure 7: Left: MATLAB code for solving the combined proportional and continuous DDE (3.9). The Chebfun function cumsummat(n,dom) computes an n-by-n spectral collocation matrix representing indefinite integration. Centre: Computed solution on an 18-point grid. Right: Geometric convergence as the discretisation size, $n$, is increased. Here the $n=30$ discretisation is used a proxy for the unknown true solution.
@@ 3.7 Other types of DDE problems
We have described above how the barycentric resampling approach can be incorporated into standard spectral collocation methods to solve the most common types of DDE problems. The approach can be further generalised to neutral-type DDEs, with a discrete operator $P(\bm{\tau},\bm{t})D$ representing $y^{\prime}(\tau(t))$, as well as higher-order problems, DDE boundary-value problems, systems of DDEs, DDE eigenvalue problems, and partial delay different equations using natural extensions of the techniques described in (Driscoll and Hale, 2016, Sections 4–6.) for ODEs. Examples of some of these are given in Section 5 below.
@@ 4 Spectral collocation for functional differential equations
Unlike time-marching methods, the approach introduced in Section 3 did not require $\tau(t)\leq t$, making it equally applicable to more general function differential equations (FDEs). As with DDEs, one must account for propagation of discontinuities caused by non-smooth initial data, as well as the potential need for an additional ‘history’ function for $t>T$ (Hale, 1971). For the sake of brevity we avoid these complications here, which can be treated in a manner similar to those described in the previous section, and instead consider two examples for which $\tau(t)\in[0,T]$.
@@
Refer to captionRefer to caption
Figure 8: Left: MATLAB code for solving the FDE (4.1) using Chebyshev spectral collocation. The only change from solving the DDE (3.1) is the first argument of barymat on line 6 and the definition of the righ hand-side vector, b. Centre: Solution using 12-point Chebyshev grid. Right: Convergence as $n$ is increased. As in the DDE case, geometric convergence to around the level of machine precision is observed.
@@
Here we consider a FDE variant of the DDE in Example 6 in which $y(t)\not<t$, namely [^8] This FDE was the original motivation for this work after a student in the author’s undergraduate class became confused when solving the much more simple ODE $y^{\prime}=-y^{2}$, $y(0)=1$ and asked if one could solve such a “crazy-wicked” equation as (4.2)!
@@
From the initial condition it follows that $y(t)<0$ and it turns out that $y(t)\in[0,1]$ for $t\in[0,1]$, meaning that no history functions or domain subdivisions are required here. Again, the fact that this is an FDE rather than a DDE has almost no impact on the discretisation, and the implementation proceeds almost identically to that of Example 6. Figure 9 shows the convergence of the Newton iteration, the computed solution, and its convergence.
Newton iteration:
iter    norm(res)        norm(dy)
@@
Refer to captionRefer to caption
Figure 9: Computed solution to the state-dependent FDE (4.2). Left: Quadratic convergence of the Newton iteration applied to the $n=12$ discretisation with initial guess $y=1$. Centre: Computed solution taking $n=12$. Right: Geometric convergence as $n$ is increased, taking the $n=30$ solution as a proxy for the unknown exact solution.
@@ 5 Chebfun & more examples
5.1 Chebfun example 1 (proportional and continuous delay):
@@
with $y(0)=e^{-1}$, with exact solution $y(t)=e^{\frac{1}{10}t-1}$. This example appears in (Wei and Chen, 2012), where it is solved using a Legendre-based spectral collocation method. The short code in Figure 10 demonstrates the ease with which this DDE can be solved in Chebfun. Here the computed solution, taking $q=\frac{1}{2}$, is returned as a length 14 chebfun and has a relative error of around $1.29\times 10^{-15}$.
1   dom = [0 20]; q = 1/2;
2   N = chebop(@(t,y) diff(y) - 1/100*(q*t-t-10)*y(q*t) - ...
@@
5   N.lbc = exp(-1);
6   y = N\0;
7   sol = chebfun(@(t) exp(t/10-1), dom);
8   err = norm(y-sol, inf)./norm(sol, inf)
Figure 10: Left: Chebfun code to solve the DDE (5.1). Right: Computed solution and relative error.
5.2 Chebfun example 2 (nonlinear, neutral type):
@@
This DDE was introduced by Jackiewicz (1984) and appears in the MATLAB example ddex5 where it solved using the built-in MATLAB routine for neutral DDEs, ddensd (MathWorks, 2012). The DDE has two consistent initial slopes, $s=2$ and $s=-W(-2e^{-2})=0.406\ldots$, where here $W$ is the Lambert W function, each corresponding to a different solution (Shampine, 2008). Chebfun code for computing both solutions is given in the listings a the top of Figure 11. When $s=2$, the DDE has an explicit solution $y=e^{\sin(2t)}$, and both the Chebfun and ddensd show good agreement with this (with relative errors of around $10^{-14}$ and $10^{-5}$, respectively), as shown in the left panel of Figure 11. For $s=-W(-2e^{-2})$ there is no known closed form solution. In the right panel of Figure 11 we see that the numerical solututions computed by Chebfun and ddesnd differ considerably, even over a relatively short time interval. This author believes the cause of this is that the DDE is highly ill-conditioned in this case, but that this is mitigated by the fully implicit (i.e., global) nature and high accuracy of the interpolation approach implemented in Chebfun. We defer further investigation to future work.
@@
Figure 11: Top: Chebfun code to solve (5.2). The Chebfun function deriv(y,t/2) evaluates $y^{\prime}$ at $t/2$. Bottom left: Agreement of Chebfun and ddesnd solutions when $s=2$. Bottom right: Disagreement of Chebfun and ddesnd solutions when $s=W(-2e^{-2})$.
5.3 Chebfun example 3 (functional, state-dependent)
Here we simply show how Example 4.2 can be implemented as a three-line code in Chebfun. The solution is depicted in Figure 6.
@@ 5.3 Chebfun example 3 (functional, state-dependent)
Figure 12: Chebfun code to solve (4.2).
5.4 Chebfun example 4 (unknown delay)
Here we revisit Example 2 from Section 3.1. However, rather than specify the proportional delay, $p$, we consider this an unknown which must be solved in order to satisfy the a two-point DDE boundary-value problem
@@ 5.4 Chebfun example 4 (unknown delay)
Figure 13: Left: Chebfun code to solve (5.3). Right: Computed solution.
5.5 Chebfun example 5 (eigenvalue problem)
As our final example in this section we consider a delay differential eigenvalue problem. [^9] The most common eigenvalue problem associated with functional or delay differential equations is that of obtaining the characteristic roots of the Jacobian at critical points to determine their stability properties. This is not what are doing here. Chebfun solves differential eigenvalue problems by computing an increasingly large Chebyshev spectral discretisation of the differential operator, as described in (Driscoll and Hale, 2016, Section 5), and looking for the ‘smoothest’ eigenvalues. With the barycentric resampling approach, this is easily adapted to the delay or functional case. Here we consider the DDE eigenvalue problem
@@ 5.5 Chebfun example 5 (eigenvalue problem)
Figure 14: Left: Chebfun code. Centre: Computed eigenvalues: Right: Eigenfunctions.
@@ 6.1 Laguerre and Hermite spectral methods
Laguerre and Hermite spectral collocation methods employ weighted-polynomial bases, $p(t)e^{-t/2}$ and $p(t)e^{-t^{2}/2}$, respectively. The approach described in Sections 3 and 4 can therefore be naturally extended to functional and delay differential defined on semi-infinite and infinite domains. For example, if $\bm{t}=[t_{1},\ldots,t_{n}]^{\top}$ is an $n$-point Gauss–Laguerre (or Gauss–Laguerre–Lobatto) grid and $\bm{w}=[w_{1},\ldots,w_{n}]^{\top}$ are the associated barycentric weights , then a functional operator $y(qt+p)$ can be discretised as
${\rm diag}(e^{-(q\bm{t}+p)/2})P(q\bm{t}+p;\bm{t}){\rm diag}(e^{\bm{t}/2})\bm{y}.$
This approach has recently been employed in (Hale et al., 2024) in the numerical investigation of the nonlinear functional differential equation
$u^{\prime}(t)+u(t)=u^{2}(\alpha t),\quad\alpha>0,\ t>0,\qquad u(0)=1.$
@@ 6.2 Fourier spectral methods
Thus far we have focussed on spectral methods based on polynomial interpolation. However, similar ideas can be applied when solving periodic DDEs using trigonometric interpolants. In particular, the degree $n$ trigonometric polynomial interpolant of an arbitrary function at $2n$ equidistant nodes, $\theta_{k}=k\pi/n$, $k=0,\ldots,2n-1$ on the interval $[0,2\pi]$ may be expressed in barycentric form as (Henrici, 1979)
@@
and a trigonometric barycentric resampling matrix constructed as in the left hand-side of Figure 2. This matrix can be combined with the trigonometric differentiation matrix (see (Trefethen, 2000, Chapter 1)) to solve DDEs on periodic domains in a similar manner to that described in Section 3. An advantage here is that the assumption of periodicity prevents the introduction of discontinuities from non-smooth initial data.
@@
Let us first consider a simple linear periodic DDE, namely
@@
with $u(t)$ $2\pi$-periodic. Code to solve this periodic problem is show in the right hand-side of Table 2. Figure 15 shows the computed solution with $n=32$ (left) and the convergence (centre) of the approach as the discretisation size is increased. In this case, we do not have a closed form expression for this solution and so take the $n=32$ solution as a proxy. The Fourier coefficients of the $n=32$ solution are shown in the rightmost panel. As in the case of the non-periodic problems, we observed geometric convergence in the number of degrees of freedom, in this case down to an accuracy of around $10^{-12}$.
@@ 6    P(isnan(P)) = 1;
1t = linspace(0, 2*pi, 2*n+1)’; t(end) = [];
2D1 = diffmat(2*n, 1, ’periodic’, [0 2*pi]);
3D2 = diffmat(2*n, 2, ’periodic’, [0 2*pi]);
@@
Table 2: Left: Matlab code for constructing the trigonometric barycentric resampling matrix for an even number of interpolation nodes t. For an odd number of points, simply replace cot with csc. Right: Matlab code for solving (6.2). Here we use diffmat from Chebfun to compute the first- and second-order trigonometric differentiation matrices. TODO: This code needs updating.
Refer to caption Refer to caption Refer to caption
Figure 15: Left: Solution using two 32-point uniform grid. Centre: Convergence Right: Fourier coefficients.
@@
A more interesting application is the use of these methods in the accurate computation of limit cycles for autonomous DDE initial value problems, which can subsequently be used in the computation of the relevant Floquet multipliers (Luzyanina and Engelborghs, 2002). For example, consider the following Lotka–Volterra predator-prey model with a delayed Holling type-II response:
@@
    For some suitable initial condition $(x_{0},y_{0})$ in the neighbourhood of the equilibrium point $(x_{e},y_{e})=(\tfrac{\alpha}{1-\alpha},(1-\tfrac{x_{e}}{K})(1+x_{e}))$, solve (6.3) for some time to obtain an approximation to the limit cycle and its unknown period, $T$.
@@
    Use the Fourier spectral collocation method with baryentric resampling to implement the delay and estimates of the limit cycle and its period from (b) as an initial guess, solve the periodic nonlinear DDE for $x$, $y$, and $T$.
@@
A similar approach is employed by (Borgioli et al., 2020) in the context of stability analysis of linear time-periodic delayed dynamical systems with piecewise-smooth coefficients. In Figure we show the result of this procedure when solving (6.3) with $K=1.4$, $\gamma=2/15$, $\delta=1$, and $s=1$, for which the period of the limit cycle is found to be 30.83847284. Checking the residual of the solution with Chebfun gives a forward error of around $6\times 10^{-14}$.
Refer to caption Refer to caption Refer to caption
Figure 16: Left: Limit cycle solution computed using 129-point grid. Centre: Limit cycle. Right: Fourier coefficients of solution.

### v2

@@ 3.6 Continuous delay
Continuous delays (sometimes called distributed delays) result in Volterra-type integro-differential equations. Driscoll (2010) presents a detailed description of a Chebyshev spectral collocation method for such equations, and we refer the reader to that work for further information. We include an example here both for completeness and to demonstrate that, since both methods are natural extensions to standard Chebyshev spectral methods for ODEs, they can be effortlessly combined to solve problems with a combination of both discrete, proportional, and continuous delay types.
@@
The code in the left panel of Figure 7 demonstrates compact MATLAB code for solving this problem, combining the resampling approach of Section 3.1 with code from (Driscoll, 2010) for discretising the Volterra operator. The computed solution is depicted in the centre panel and geometric convergence as the discretisation size is increased is demonstrated on the right.
1K = @(s,t) exp(-(t-s).^2);
2n = 14; dom = [0 1];
3t = chebpts(n, dom);
4D = diffmat(n, dom);
5P = Barymat(t/2, t);
6V = K(t, t’).*cumsummat(n, dom);
7A = D + P/2 - V ; rhs = 0*t;
8A(1,:) = eye(1,n); rhs(1) = 1;
@@
Figure 7: Left: MATLAB code for solving the combined proportional and continuous DDE (3.19). The Chebfun function cumsummat(n,dom) computes an n-by-n spectral collocation matrix representing indefinite integration. Note that here we use the Chebfun implementations of diffmat and barymat, which are hardcoded for Chebyshev points. Centre: Computed solution on an 14-point grid. Right: Geometric convergence as the discretisation size, $n$, is increased. Here the $n=30$ discretisation is used a proxy for the unknown true solution.
@@ 3.7 Other types of DDE problems
We have described above how the barycentric resampling approach can be incorporated into standard spectral collocation methods to solve the most common types of DDE problems. The approach can be further generalised to neutral-type DDEs, with a discrete operator $P(\bm{\tau},\bm{t})D$ representing $y^{\prime}(\tau(t))$, as well as higher-order problems, DDE-BVPs, systems of DDEs, DDE-EVP, and partial delay differential equations using natural extensions of the techniques described in (Driscoll and Hale, 2016, Sections 4–6.) for ODEs. Examples of some of these are given in Section 5 below.
@@ 4 Spectral collocation for functional differential equations
Unlike time-marching methods, the approach introduced in Section 3 did not require $\tau(t)\leq t$, making it applicable to more general functional differential equations (FDEs). As with DDEs, one must account for propagation of discontinuities caused by non-smooth initial data, as well as the potential need for an additional ‘history’ function for $t>T$ (Hale, 1977). For the sake of brevity we avoid these complications here, which can be treated in a manner similar to those described in the previous section, and instead consider two examples for which $\tau(t)\in[0,T]$.
@@
Figure 8: Left: MATLAB code for solving the FDE (4.1) using Chebyshev spectral collocation. The only change from solving the DDE (3.1) is the first argument of Barymat on line 5 and the definition of the right-hand side vector, b. Centre: Solution using 12-point Chebyshev grid. Right: Convergence as $n$ is increased. As in the DDE case, geometric convergence to around the level of machine precision is observed.
@@
Here we consider an FDE variant of the DDE in Example 6 in which $y(t)\not<t$, namely [^8] This FDE was the original motivation for this work after a student in the author’s undergraduate class became confused when solving the much more simple ODE $y^{\prime}=-y^{2}$, $y(0)=1$ and asked how one might solve such an equation as (4.2). [^9] A DDE closely related to (4.2) is considered in some detail by Eder (1984).
@@
Here we are lucky in that $y(t)\in[0,1]$ for $t\in[0,1]$, meaning that no history functions or domain subdivisions are required. Again, the fact that this is an FDE rather than a DDE has almost no impact on the discretisation, and the implementation proceeds almost identically to that of Example 6. Figure 9 shows convergence of the Newton iteration, the computed solution, and its convergence.
$k$ Residual norm Update norm
@@
Figure 9: Computed solution to the state-dependent FDE (4.2). Left: Quadratic convergence of the Newton iteration applied to the $n=12$ discretisation with initial guess $y=1$. Centre: Computed solution taking $n=12$. Right: Geometric convergence as $n$ is increased, taking the $n=30$ solution as a proxy for the unknown true solution.
@@ 5 Chebfun & more examples
5.1 Chebfun example 1. (proportional and continuous delay)
@@
for $t\in[0,20]$ with $y(0)=e^{-1}$, with exact solution $y(t)=e^{\frac{1}{10}t-1}$. This example appears in (Wei and Chen, 2012), where it is solved using a Legendre-based spectral collocation method. The short code in Figure 10 demonstrates the ease with which this DDE can be solved in Chebfun. Here the computed solution, taking $q=\frac{1}{2}$, is returned as a length 14 chebfun and has a relative error of around $1.5\times 10^{-15}$.
1dom = [0 20]; q = 1/2;
2N = chebop(@(t,y) diff(y) - 1/100*(q*t-t-10)*y(q*t) - ...
@@
5N.lbc = exp(-1);
6y = N\0;
7sol = chebfun(@(t) exp(t/10-1), dom);
8err = norm(y-sol, inf)./norm(sol, inf)
Figure 10: Left: Chebfun code to solve the continuous-delay DDE (5.1). Right: Computed solution (dots) and relative error.
5.2 Chebfun example 2. (nonlinear, neutral type)
@@
This DDE was introduced by Jackiewicz (1984) and appears in the MATLAB example ddex5 where it is solved using the built-in MATLAB routine for neutral DDEs, ddensd (MathWorks, ). By substituting $y(0)=1$ and $y^{\prime}(0)=s$ to (5.2), one can show that the DDE has two consistent initial slopes (i.e., choices of $y^{\prime}(0)$ so that (5.2) is satisfied at $t=0$), namely $s=2$ and $s=-W(-2e^{-2})=0.406\ldots$, where here $W$ is the Lambert W function, each corresponding to a different solution (Shampine, 2008). Chebfun code for computing both solutions is given at the top of Figure 11. When $s=2$, the DDE has an explicit solution $y=e^{\sin(2t)}$, and both the Chebfun and ddensd show good agreement with this (with relative errors of around $10^{-14}$ and $10^{-5}$, respectively), as shown in the left panel of Figure 11. For $s=-W(-2e^{-2})$ there is no known closed form solution. In the right panel of Figure 11 we see that the numerical solutions computed by Chebfun and ddesnd differ considerably, even over a relatively short time interval. This author believes the cause of this is that the DDE is highly ill-conditioned in this case, but that this is mitigated by the fully implicit (i.e., global) nature and high accuracy of the interpolation approach implemented in Chebfun. We defer further investigation to future work.
@@
Figure 11: Top: Chebfun code to solve the neutral-type DDE (5.2). The Chebfun function deriv(y,t/2) evaluates $y^{\prime}$ at $t/2$. Bottom left: Agreement of Chebfun and ddesnd solutions when $s=2$. Bottom right: Disagreement of Chebfun and ddesnd solutions when $s=W(-2e^{-2})$.
5.3 Chebfun example 3. (functional, state-dependent)
Here we simply show how Example 9 can be implemented as a three-line code in Chebfun. The solution is depicted in Figure 6.
@@
5.4 Chebfun example 4. (unknown delay)
Here we revisit Example 2 from Section 3.1. However, rather than specify the proportional delay, $p$, we consider this an unknown which must be solved in order to satisfy the two-point DDE boundary-value problem
@@
Figure 12: Left: Chebfun code to solve the unknown delay problem (5.3). Right: Computed solution.
5.5 Chebfun example 5. (eigenvalue problem)
As our final example in this section we consider a delay differential eigenvalue problem (DDE-EVP). [^10] The most common eigenvalue problem associated with functional or delay differential equations is that of obtaining the characteristic roots of the Jacobian at critical points to determine their stability properties. This is not what we are doing here. Chebfun solves differential eigenvalue problems by computing an increasingly large Chebyshev spectral discretisation of the differential operator and looking for the ‘smoothest’ eigenvalues, as described in (Driscoll and Hale, 2016, Section 5). With the barycentric resampling approach, this is easily adapted to the delay or functional case. Here we consider the second-order DDE-EVP
@@
Figure 13: Left: Chebfun code to solve to solve the DDE-EVP (5.4). Centre: Computed eigenvalues: Right: Computed eigenfunctions.
@@ 6.1 Laguerre and Hermite spectral methods
Laguerre and Hermite spectral collocation methods, which employ weighted-polynomial bases, $\psi_{L}(t)=e^{-bt/2}p(t)$ and $\psi_{H}(t)=e^{-bt^{2}/2}p(t)$, $b>0$, have been widely used in solving ODEs and PDEs on semi-infinite and infinite domains, respectively (Boyd, 2001, Chapter 17). The approach described in Sections 3 and 4 above can be naturally extended to solve functional and delay differential equations using these methods as follows. Consider, for example, the Laguerre case, where the weighted-polynomial interpolant may be represented as $\psi_{L}(t)=e^{-bt/2}{P(t;\bm{t})(e^{b\bm{t}/2}\cdot\bm{\psi})}$, where $\bm{t}=[t_{1},\ldots,t_{n}]^{\top}$ is an $n$-point Gauss–Laguerre (or Gauss–Laguerre–Lobatto) grid, $P(t;\bm{t})$ is the Lagrange interpolation quasimatrix (2.2), and $\bm{\psi}=\psi_{L}(\bm{t})$. It follows that $\psi_{L}(\tau)=e^{-b\tau/2}{P(\tau;\bm{t})(e^{b\bm{t}/2}\cdot\bm{\psi})}$ and the interpolation operator can be discretised as $y(\bm{\tau})\approx P_{L}(\bm{\tau};\bm{t})y(\bm{t}),$ where
$P_{L}(\bm{\tau};\bm{t})={\rm diag}(e^{-b\bm{\tau}/2})P(\bm{\tau};\bm{t}){\rm diag}(e^{b\bm{t}/2}),$ (6.1)
and $P(\bm{\tau};\bm{t})$ is precisely the barycentric resampling matrix from Section 2.1. Corresponding differentiation matrices can be constructed using, for example, DMSUITE (Weideman and Reddy, 2000), and combined with the above as in previous sections. This approach has recently been employed by Hale et al. (2024) for the numerical investigation of the nonlinear functional differential equation
$u^{\prime}(t)+u(t)=u^{2}(\alpha t),\quad\alpha>0,\ t>0,\qquad u(0)=1,$ (6.2)
which yields the nonlinear system
$(D+I)\bm{u}-P_{L}(\alpha\bm{t};\bm{t})(\bm{u}\cdot\bm{u})=0.$ (6.3)
@@ 6.2 Fourier spectral methods
Thus far we have focussed on spectral methods based on polynomial interpolation. However, similar ideas can be applied when solving PDDEs using trigonometric interpolants. In particular, the degree $n$ trigonometric polynomial interpolant of an arbitrary function at $2n$ equidistant nodes, $\theta_{k}=k\pi/n$, $k=0,\ldots,2n-1$ on the interval $[0,2\pi]$ may be expressed in barycentric form as (Henrici, 1979)
@@
and a trigonometric barycentric resampling matrix constructed as in the left-hand side of Figure 2. This matrix can be combined with the trigonometric differentiation matrix—see (Trefethen, 2000, Chapter 1)—to solve DDEs on periodic domains in a similar manner to that described in Section 3. An advantage here is that the assumption of periodicity prevents the introduction of discontinuities from non-smooth initial data.
@@
Let us first consider a simple linear PDDE, namely
@@
with $u(t)$ $2\pi$-periodic. Code to solve this periodic problem is show in the right-hand side of Table 2. Figure 14 shows the computed solution with $n=32$ (left) and the convergence (centre) of the approach as the discretisation size is increased. In this case, we do not have a closed form expression for this solution and so take the $n=32$ solution as a proxy. The Fourier coefficients of the $n=32$ solution are shown in the rightmost panel. As in the case of the non-periodic problems, we observe geometric convergence in the number of degrees of freedom, in this case down to an accuracy of around $10^{-12}$.
@@ 6 P(isnan(P)) = 1;
1t = linspace(0, 2*pi, n+1)’; t(end) = [];
2D1 = diffmat(n, 1, ’periodic’, [0 2*pi]);
3D2 = diffmat(n, 2, ’periodic’, [0 2*pi]);
@@
Table 2: Left: MATLAB code for constructing the trigonometric barycentric resampling matrix for an even number of interpolation nodes t. For an odd number of points, simply replace cot with csc. Right: MATLAB code for solving (6.5). Here we use diffmat from Chebfun to compute the first- and second-order trigonometric differentiation matrices, D1 and D2.
Figure 14: Left: Solution to the periodic DDE (6.5) using Fourier collocation on a 32-point uniform grid. Centre: Geometric convergence as $n$ is increased. Right: Fourier coefficients of the computed $n=32$ solution, exhibiting geometric decay.
@@
A more interesting application is the use of these methods in the accurate computation of limit cycles for autonomous DDEs, which can subsequently be used in the computation of the relevant Floquet multipliers (Luzyanina and Engelborghs, 2002). For example, consider the following Lotka–Volterra predator-prey model with a delayed Holling type-II response:
@@
    For some suitable initial condition $(x_{0},y_{0})$ in the neighbourhood of the equilibrium point $(x_{e},y_{e})=(\tfrac{\alpha}{1-\alpha},(1-\tfrac{x_{e}}{K})(1+x_{e}))$, solve (6.6) for some time using an appropriate time-integration routine (such as dde23 in MATLAB) to obtain an approximation to the limit cycle and its unknown period, $T$.
@@
    Use the Fourier spectral collocation method with barycentric resampling to implement the delay and estimates of the limit cycle and its period from (b) as an initial guess, and solve the periodic nonlinear DDE for $x$, $y$, and $T$.
@@
A similar approach is employed by (Borgioli et al., 2020) in the context of stability analysis of linear time-periodic delayed dynamical systems with piecewise-smooth coefficients. In Figure 15 we show the result of this procedure when solving (6.6) with $K=7/5$, $\gamma=2/15$, $\delta=1$, and $s=1$, for which the period of the limit cycle is found to be $T=30.83847284$. Checking the residual of the solution with Chebfun gives a forward error of around $6\times 10^{-14}$. The solution is depsicted in Figure 15.
Figure 15: Left: Limit cycle solution of the delayed Lotka–Volterra DDE (6.6), computed using a 129-point grid. Centre: Initial trajectory computed by dde23 (dashed) and limit cycle computed by the Fourier pseudospectral method (solid). Right: Fourier coefficients of the computed limit cycle solution, demonstrating geometric decay.